A shelf can hold 7 trophies. How many ways can the trophies be arranged if there are 10 trophies available?
Answer by stanbon(75887)

A shelf can hold 7 trophies. How many ways can the trophies be arranged if there are 10 trophies available?
-----
Ans: 10P7 = 10!/(10-7)! = 10*9*8*7*6*5*4 = 604800 ways
-----------
Cheers,
Stan H.
----------